OpenStudy (anonymous):

help please ??find any intercepts y=2x- √x²+1 thanks (: !

OpenStudy (anonymous):

graph it, thats the easiest

OpenStudy (anonymous):

other than that you could do: set the \(y\) equal to \(0\) and solve for \(x\) to find the x-intercept:\[2x-\sqrt{x}^2+1=0\]simplify the radical:\[2x-x+1=0\]combine like terms:\[x+1=0\]subtract 1 to isolate \(x\):\[x=-1\]^so that's the x-intercept now to find the y-intercept, set all the \(x\)'s equal to \(0\):\[2(0)-\sqrt{0}^2+1=y\]simplify the radical:\[2(0)-0+1=y\]multiply:\[0-0+1=y\]combine like terms:\[1=y\] \[\text{the x and y intercepts are (-1,0) and (0,1), respectively.}\] hope this helps! :) @Ephilo

OpenStudy (anonymous):

@yummydum yes ! a lot !!

OpenStudy (anonymous):

\[\huge\ddot\smile\]

OpenStudy (anonymous):

@yummydum im stuck another one can you help me with it too? please

OpenStudy (anonymous):

sure, post another question

OpenStudy (anonymous):

i kind of get this one but i'm lost in the middle Find the points of intersection of the graphs of the equations: x²+y²=25 2x+y=10

OpenStudy (anonymous):

@yummydum ^^

OpenStudy (anonymous):

for this one you'll have to graph graph them both and see where they intersect

OpenStudy (anonymous):

okay thanks (: